# VisMem: Latent Vision Memory Unlocks Potential of Vision-Language Models

### v1

@@ Abstract
Despite the remarkable success of Vision-Language Models (VLMs), their performance on a range of complex visual tasks is often hindered by a “visual processing bottleneck”: a propensity to lose grounding in visual evidence and exhibit a deficit in contextualized visual experience during prolonged generation. Drawing inspiration from human cognitive memory theory, which distinguishes short-term visually-dominant memory and long-term semantically-dominant memory, we propose VisMem, a cognitively-aligned framework that equips VLMs with dynamic latent vision memories, a short-term module for fine-grained perceptual retention and a long-term module for abstract semantic consolidation. These memories are seamlessly invoked during inference, allowing VLMs to maintain both perceptual fidelity and semantic consistency across thinking and generation. Extensive experiments across diverse visual benchmarks for understanding, reasoning, and generation reveal that VisMem delivers a significant average performance boost of 11.8% relative to the vanilla model and outperforms all counterparts, establishing a new paradigm for latent-space memory enhancement. The code will be available: https://github.com/YU-deep/VisMem.git.
@@ 1 Introduction
Visual-Language Models (VLMs) have demonstrated impressive capabilities in visual understanding, reasoning and generation [li2024survey, su2025thinking]. Latest flagship models, both closed-sourced [gpt5, gemini, anthropic2025claude] and open-sourced [bai2025qwen2, team2025kimi, an2025llava, wang2025internvl3, vteam2025glm45v], represent a significant leap towards a general-purpose intelligent model that can both perceive and think about the visual world. Despite their success, VLMs still face significant inherent challenges when tackling complicated tasks that require advanced visual abilities, such as fine-grained perception, multi-step reasoning, or maintaining fidelity over long generative sequences [ghosh2024exploring, huang2025visfactor]. A fundamental limitation stems from the pervasive propensity, exhibited during deep autoregressive decoding, toward a deficit in visual memory, which prioritizes accumulated textual context over the initial visual evidence and lacks visual semantic knowledge [zhou2024rethinking, sun2025mitigating]. It manifests as a “visual processing bottleneck” that impairs performance in fine-grained visual understanding, efficient reasoning, and robust generation.
Prior efforts to overcome this limitation have explored several distinct strategic axes, which can be primarily categorized into four paradigms, as illustrated in Fig. 1. One intuitive paradigm is the (a) direct training paradigm, which optimizes model parameters via fine-tuning or reinforcement learning [liu2025visual, shen2025vlm, huang2025vision, wang2025perception]. This relatively brute-force approach often sacrifices generalization for task-specific performance, leading to catastrophic forgetting. Another axis concerns the representation space of the intervention, (b) image-level paradigm, operating in the pixel space by explicitly synthesizing new visual inputs, which offers image-level thinking but at a prohibitive computational cost [zheng2025deepeyes, fan2025grit, hu2024visual, su2025pixel, li2025imagine, su2025openthinkimg]. Conversely, (c) token-level paradigm constrains operations to visual tokens, which is more efficient but fundamentally non-generative, limiting the model to merely re-surfacing what it has already encoded [lei2024scaffolding, gao2025interleaved, chen2025mint, yu2025introducing]. Recently, a promising direction lies in the (d) latent space paradigm, which introduces continuous latent contexts in the sequential inference process. Unfortunately, existing latent space methods either rely solely on the language space [hao2024training, xu2025softcot, shen2025codi, li2025seek, zhang2025memgen] or require auxiliary visual data [yang2025machine], limiting their application in VLMs.
@@
To overcome this problem, we resort to cognitive psychology, specifically the Dennis Norris Theory [norris2017short]:
@@ 2.1 Visual Capacities Enhancement
As demonstrated in Fig. 1, existing methods to alleviate “visual processing bottleneck” of VLMs broadly fall into four main categories: (a) direct training paradigm, which directly optimizes model parameters for target visual tasks, as in SFT, Visual-RFT [liu2025visual], VLM-R1 [shen2025vlm], Vision-R1 [huang2025vision], and PAPO [wang2025perception]. Nonetheless, these methods suffer from catastrophic forgetting, specifically manifested as the degradation of general capabilities and over-specialization in specific visual cognition tasks [zhou2023learning, yu2024boosting]; (b) image-level paradigm, which either leverages bounding boxes to denote visual evidence, represented by methods as Visual CoT [shao2024visual], DeepEyes [zheng2025deepeyes], SpatialVTS [liang2025enhancing], VGR [wang2025vgr], and GRIT [fan2025grit], or externally generate the iterative visual inputs via predefined tools, as seen in Sketchpad [hu2024visual], VPRL [xu2025visual], PyVision [zhao2025pyvision], OpenAI o3 [openai2025o3], PixelReasoner [su2025pixel], MVoT [li2025imagine], and OpenThinkImg [su2025openthinkimg]. Nevertheless, modifying visual inputs incurs extremely high computational costs, accompanied by high latency and reliance on external tools and concretized images; (c) token-level paradigm, which select original representations and cannot modify visual evidences, thus restricted by insufficiently refined information and suboptimal selection strategies, as in ICoT [gao2025interleaved], MINT-CoT [chen2025mint], SCAFFOLD [lei2024scaffolding], LLaVA-AURORA [bigverdi2025perception], VPT [yu2025introducing], Chameleon [team2024chameleon], (d) latent space paradigm, which employs latent states to optimize autoregressive generation, but its focus remains on pure language models, e.g., Coconut [hao2024training], MemGen [zhang2025memgen], LatentSeek [li2025seek], SoftCoT [xu2025softcot], CODI [shen2025codi]. Although Mirage [yang2025machine] attempts to construct a latent vision space, requiring substantial manually labeled images. Our VisMem also belongs to this paradigm, but differs from existing methods by integrating latent vision memory within generation processes, characterized by a short and long memory system.
@@ 2.2 Memory Empowerment
Another mechanism closely tied to our approach involves endowing models with memory functionality. One intuitive strategy entails directly optimize models on prior trajectories, exemplified by [shen2024memory, zhang2025g, fu2025agentrefine], or to store them into the external memory repositories [tack2024online, wang2024wise]. Besides, some models inject persistently stored, retrieval-augmented knowledge from external environments, such as Expel [zhao2024expel] and MemoryBank [zhong2024memorybank], others, such as SkillWeaver [zheng2025skillweaver] and Alita [qiu2025alita], distill prior knowledge as reusable tools. Currently, latent memory, as an implicit memory representation with better cross-domain generalization, efficiently encodes deep semantic associations, including M+ [wang2025m] and MemGen [zhang2025memgen]. Nevertheless, these memory paradigms fail to ideally accommodate visual information, which manifests as a continuous, high-dimensional perceptual input. Consequently, the exploration of efficient visual memory mechanisms remains a largely uncharted territory. Thus, we propose a more human-aligned latent vision memory paradigm.
@@ 3.1 Preliminary
Motivation. Building on the Dennis Norris Theory [norris2017short], which aligns with contemporary models of human memory, the coordinated operation of short- and long-term visual memories surmounts the “visual processing bottleneck”. Short-term latent visual memory maintains fine-grained detail for immediate use and is thus visually dominant; by contrast, long-term latent visual memory abstracts across experiences to enable flexible reuse and is therefore semantically dominant. Taking the task illustrated in Fig. 2 as a case in point, “find the classic Lay’s on the shelf” entails the deployment of short-term vision memory, retaining visual details for immediate perceptual demands, while “get in the promotion” triggers generalized semantic knowledge about the “promotion label” acquired from historical scenarios, which is grounded in long-term latent memory, to facilitate the comprehension of the task-based sight. Existing paradigms for enhancing visual capabilities fail to adequately consider vision memory, thus, our VisMem proposes a latent memory method to bridge this gap. More theoretical foundations are in Appendix 6.
@@ 3.2 Memory Invocation
As illustrated in Fig. 2, our latent vision memory invocation strategy largely aligns with the standard generation pipeline of VLMs, thereby preserving their robust fundamental visual capabilities. Typically, VLMs generate rationales and answers; however, such pure text sequences lack the granularity to capture fine-grained visual perceptions and semantics, which poses challenges to accurate visual understanding, reasoning, and generation. This limitation arises because during inference, VLMs tend to prioritize accumulated textual context over visual evidence, a phenomenon particularly pronounced in long sequences [ghosh2024exploring, yu2025visual, yin2025clearsight, huang2025visfactor]. To address this, we extend the vocabulary $\mathcal{V}$ of VLMs by incorporating four additional memory-operation tokens, resulting in $\mathcal{V}^{\prime}=\mathcal{V}\cup\left\{\texttt{<}m^{s}_{I}\texttt{>},\texttt{<}m^{s}_{E}\texttt{>},\texttt{<}m^{l}_{I}\texttt{>},\texttt{<}m^{l}_{E}\texttt{>}\right\}$. Here, $\texttt{<}m_{I}\texttt{>}$ and $\texttt{<}m_{E}\texttt{>}$ form paired invocation and end tokens, where the superscripts $s$ and $l$ denote short- or long-term memory, respectively. Specifically, we register these as indivisible special tokens in the tokenizer and enlarge the embedding matrix from $\mathbb{R}^{\mid\mathcal{V}\mid\times d}$ to $\mathbb{R}^{\left(\mid\mathcal{V}\mid+4\right)\times d}$, where $d$ is the dimension of the model. Furthermore, we initialize the embeddings of the invocation tokens ($\texttt{<}m^{s}_{I}\texttt{>}$ and $\texttt{<}m^{l}_{I}\texttt{>}$) using the embedding vector of a delimiter token with small perturbations, and update these embeddings during training to facilitate faster convergence. The two end tokens ($\texttt{<}m^{s}_{E}\texttt{>}$ and $\texttt{<}m^{l}_{E}\texttt{>}$) are treated as structural markers; they are initialized analogously with a lower learning rate. In practice, we also employ constrained decoding to encourage well-formed invocation-end pairs.
@@ 3.3 Memory Formation
Latent Memory Former. Distinct from many existing paradigms [shen2025vlm, huang2025vision, yang2025machine], we internalize the latent vision memory into lightweight formers, preserving the general abilities of base VLMs and ensuring the compatibility of our paradigm. We initialize two lightweight LoRA adapters, which are respectively designated as the short-term memory former $\mathcal{F}_{s}$ and long-term memory former $\mathcal{F}_{l}$, attached to the vision encoder and the final language model of the VLM, without directly tampering with the core parameters. More precisely, we first append the generated memory query $\mathbf{Q}$ along with a set of learnable memory tokens after the corresponding target token sequence X. Then we process it by short-term or long-term memory former, which contextualizes and embeds the latent memory information:
@@ 3.4 Training Recipe
We design a two-stage training procedure based on GRPO [shao2024deepseekmath], whose optimization objectives are to optimize the effective formation and invocation of latent memory. The first stage enhances the utility of memory, while the second stage maximizes the reward of each invocation, thereby accelerating the convergence of different components steadily. More detailed algorithms and implementations are present in Appendix 7.2 and 8.3.
@@ 4 Experiments
Star Vet Bench Vista -V -Math Bench Trust
Vanilla [bai2025qwen2] 62.6 66.0 54.0 55.4 57.4 59.3 56.0 43.5 23.8 18.9 35.6 52.3 64.8 55.4 57.7 51.0
SFT 64.7 67.5 56.8 54.5 58.7 60.3 57.7 46.1 25.9 22.8 38.1 53.6 67.0 59.1 59.9 52.8
Visual-RFT [liu2025visual] 65.6 70.5 59.1 58.0 62.9 63.6 62.4 51.7 33.4 26.5 45.2 55.8 70.7 63.2 63.2 57.4
VLM-R1 [shen2025vlm] 66.3 73.0 59.4 60.6 63.8 64.6 63.4 53.0 39.5 34.6 47.6 54.2 69.9 61.7 61.9 58.3
Vision-R1 [huang2025vision] 67.1 71.7 60.2 60.8 64.0 65.0 63.2 53.9 42.5 38.7 49.8 56.4 72.6 63.6 64.2 59.7
PAPO [wang2025perception] 64.2 69.8 57.9 53.3 56.7 60.4 61.2 52.5 34.1 34.8 45.7 50.3 67.7 56.5 58.2 55.0
Sketchpad [hu2024visual] 62.1 64.5 57.0 54.9 52.8 58.3 57.9 47.4 27.5 24.6 39.4 52.1 66.2 57.2 58.5 52.1
GRIT [fan2025grit] 65.8 67.8 57.9 52.5 51.0 59.0 59.4 51.6 30.8 22.4 43.6 53.7 67.3 60.1 60.4 54.2
PixelReasoner [su2025pixel] 65.3 67.1 58.7 56.8 60.5 61.7 58.9 49.3 32.0 25.9 43.0 55.9 69.9 61.5 62.4 55.6
DeepEyes [zheng2025deepeyes] 66.4 70.5 60.3 60.4 63.0 64.1 60.3 49.1 34.6 31.5 43.9 57.4 72.6 64.6 64.9 57.6
OpenThinkImg [su2025openthinkimg] 66.0 71.6 60.8 59.2 61.7 63.9 61.4 52.8 37.8 28.0 45.8 54.9 74.0 64.3 64.4 58.0
Scaffold [lei2024scaffolding] 63.9 67.0 58.5 52.5 52.9 59.0 58.1 51.0 29.3 21.0 41.9 54.8 68.5 60.6 61.3 53.9
ICoT [gao2025interleaved] 65.6 67.9 60.5 54.3 57.0 61.1 58.6 49.8 42.9 30.8 46.0 57.0 69.1 62.0 62.7 56.5
MINT-CoT [chen2025mint] 66.2 69.5 57.3 55.4 58.9 61.5 57.7 51.5 47.3 39.2 48.9 56.7 71.4 60.8 63.0 57.7
VPT [yu2025introducing] 64.2 70.8 59.0 58.6 63.5 63.2 59.1 53.0 37.6 34.7 46.1 52.3 64.7 61.4 59.5 56.6
Mirage [yang2025machine] 64.5 71.8 56.1 56.3 59.0 61.5 59.4 50.6 36.1 35.4 45.4 50.9 66.1 60.3 59.1 55.5

### v2

@@ Abstract
Despite the remarkable success of Vision-Language Models (VLMs), their performance on a range of complex visual tasks is often hindered by a “visual processing bottleneck”: a propensity to lose grounding in visual evidence and exhibit a deficit in contextualized visual experience during prolonged generation. Drawing inspiration from human cognitive memory theory, which distinguishes short-term visually-dominant memory and long-term semantically-dominant memory, we propose VisMem, a cognitively-aligned framework that equips VLMs with dynamic latent vision memories, a short-term module for fine-grained perceptual retention and a long-term module for abstract semantic consolidation. These memories are seamlessly invoked during inference, allowing VLMs to maintain both perceptual fidelity and semantic consistency across thinking and generation. Extensive experiments across diverse visual benchmarks for understanding, reasoning, and generation reveal that VisMem delivers a significant average performance boost of 11.0% relative to the vanilla model and outperforms all counterparts, establishing a new paradigm for latent-space memory enhancement. The code will be available: https://github.com/YU-deep/VisMem.git.
@@ 1 Introduction
Visual-Language Models (VLMs) have demonstrated impressive capabilities in visual understanding, reasoning and generation [31, 50]. Latest flagship models, both closed-sourced [39, 11, 2] and open-sourced [4, 56, 1, 63, 55], represent a significant leap towards a general-purpose intelligent model that can both perceive and think about the visual world. Despite their success, VLMs still face significant inherent challenges when tackling complicated tasks that require advanced visual abilities, such as fine-grained perception, multi-step reasoning, or maintaining fidelity over long generative sequences [17, 25]. A fundamental limitation stems from the pervasive propensity, exhibited during deep autoregressive decoding, toward a deficit in visual memory, which prioritizes accumulated textual context over the initial visual evidence and lacks visual semantic knowledge [90, 52]. It manifests as a “visual processing bottleneck” that impairs performance in fine-grained visual understanding, efficient reasoning, and robust generation.
Prior efforts to overcome this limitation have explored several distinct strategic axes, which can be primarily categorized into four paradigms, as illustrated in Fig. 1. One intuitive paradigm is the (a) direct training paradigm, which optimizes model parameters via fine-tuning or reinforcement learning [35, 44, 26, 66]. This relatively brute-force approach often sacrifices generalization for task-specific performance, leading to catastrophic forgetting. Another axis concerns the representation space of the intervention, (b) image-level paradigm, operating in the pixel space by explicitly synthesizing new visual inputs, which offers image-level thinking but at a prohibitive computational cost [87, 13, 24, 48, 29, 49]. Conversely, (c) token-level paradigm constrains operations to visual tokens, which is more efficient but fundamentally non-generative, limiting the model to merely re-surfacing what it has already encoded [28, 16, 9, 75]. Recently, a promising direction lies in the (d) latent space paradigm, which introduces continuous latent contexts in the sequential inference process. Unfortunately, existing latent space methods either rely solely on the language space [21, 69, 47, 30, 81] or require auxiliary visual data [70], limiting their application in VLMs.
@@
To overcome this problem, we resort to cognitive psychology, specifically the Dennis Norris Theory [38]:
@@ 2.1 Visual Capacities Enhancement
As demonstrated in Fig. 1, existing methods to alleviate “visual processing bottleneck” of VLMs broadly fall into four main categories: (a) direct training paradigm, which directly optimizes model parameters for target visual tasks, as in SFT, Visual-RFT [35], VLM-R1 [44], Vision-R1 [26], and PAPO [66]. Nonetheless, these methods suffer from catastrophic forgetting, specifically manifested as the degradation of general capabilities and over-specialization in specific visual cognition tasks [89, 74]; (b) image-level paradigm, which either leverages bounding boxes to denote visual evidence, represented by methods as Visual CoT [42], DeepEyes [87], SpatialVTS [33], VGR [58], and GRIT [13], or externally generate the iterative visual inputs via predefined tools, as seen in Sketchpad [24], VPRL [68], PyVision [85], OpenAI o3 [40], PixelReasoner [48], MVoT [29], and OpenThinkImg [49]. Nevertheless, modifying visual inputs incurs extremely high computational costs, accompanied by high latency and reliance on external tools and concretized images; (c) token-level paradigm, which select original representations and cannot modify visual evidences, thus restricted by insufficiently refined information and suboptimal selection strategies, as in ICoT [16], MINT-CoT [9], SCAFFOLD [28], LLaVA-AURORA [6], VPT [75], Chameleon [54], (d) latent space paradigm, which employs latent states to optimize autoregressive generation, but its focus remains on pure language models, e.g., Coconut [21], MemGen [81], LatentSeek [30], SoftCoT [69], CODI [47]. Although Mirage [70] attempts to construct a latent vision space, requiring substantial manually labeled images. Our VisMem also belongs to this paradigm, but differs from existing methods by integrating latent vision memory within generation processes, characterized by a short and long memory system.
@@ 2.2 Memory Empowerment
Another mechanism closely tied to our approach involves endowing models with memory functionality. One intuitive strategy entails directly optimize models on prior trajectories, exemplified by [45, 80, 14], or to store them into the external memory repositories [53, 62]. Besides, some models inject persistently stored, retrieval-augmented knowledge from external environments, such as Expel [83] and MemoryBank [88], others, such as SkillWeaver [86] and Alita [41], distill prior knowledge as reusable tools. Currently, latent memory, as an implicit memory representation with better cross-domain generalization, efficiently encodes deep semantic associations, including M+ [64] and MemGen [81]. Nevertheless, these memory paradigms fail to ideally accommodate visual information, which manifests as a continuous, high-dimensional perceptual input. Consequently, the exploration of efficient visual memory mechanisms remains a largely uncharted territory. Thus, we propose a more human-aligned latent vision memory paradigm.
@@ 3.1 Preliminary
Motivation. Building on the Dennis Norris Theory [38], which aligns with contemporary models of human memory, the coordinated operation of short- and long-term visual memories surmounts the “visual processing bottleneck”. Short-term latent visual memory maintains fine-grained detail for immediate use and is thus visually dominant; by contrast, long-term latent visual memory abstracts across experiences to enable flexible reuse and is therefore semantically dominant. Taking the task illustrated in Fig. 2 as a case in point, “find the classic Lay’s on the shelf” entails the deployment of short-term vision memory, retaining visual details for immediate perceptual demands, while “get in the promotion” triggers generalized semantic knowledge about the “promotion label” acquired from historical scenarios, which is grounded in long-term latent memory, to facilitate the comprehension of the task-based sight. Existing paradigms for enhancing visual capabilities fail to adequately consider vision memory, thus, our VisMem proposes a latent memory method to bridge this gap. More theoretical foundations are in Appendix 6.
@@ 3.2 Memory Invocation
As illustrated in Fig. 2, our latent vision memory invocation strategy largely aligns with the standard generation pipeline of VLMs, thereby preserving their robust fundamental visual capabilities. Typically, VLMs generate rationales and answers; however, such pure text sequences lack the granularity to capture fine-grained visual perceptions and semantics, which poses challenges to accurate visual understanding, reasoning, and generation. This limitation arises because during inference, VLMs tend to prioritize accumulated textual context over visual evidence, a phenomenon particularly pronounced in long sequences [17, 78, 72, 25]. To address this, we extend the vocabulary $\mathcal{V}$ of VLMs by incorporating four additional memory-operation tokens, resulting in $\mathcal{V}^{\prime}=\mathcal{V}\cup\left\{\texttt{<}m^{s}_{I}\texttt{>},\texttt{<}m^{s}_{E}\texttt{>},\texttt{<}m^{l}_{I}\texttt{>},\texttt{<}m^{l}_{E}\texttt{>}\right\}$. Here, $\texttt{<}m_{I}\texttt{>}$ and $\texttt{<}m_{E}\texttt{>}$ form paired invocation and end tokens, where the superscripts $s$ and $l$ denote short- or long-term memory, respectively. Specifically, we register these as indivisible special tokens in the tokenizer and enlarge the embedding matrix from $\mathbb{R}^{\mid\mathcal{V}\mid\times d}$ to $\mathbb{R}^{\left(\mid\mathcal{V}\mid+4\right)\times d}$, where $d$ is the dimension of the model. Furthermore, we initialize the embeddings of the invocation tokens ($\texttt{<}m^{s}_{I}\texttt{>}$ and $\texttt{<}m^{l}_{I}\texttt{>}$) using the embedding vector of a delimiter token with small perturbations, and update these embeddings during training to facilitate faster convergence. The two end tokens ($\texttt{<}m^{s}_{E}\texttt{>}$ and $\texttt{<}m^{l}_{E}\texttt{>}$) are treated as structural markers; they are initialized analogously with a lower learning rate. In practice, we also employ constrained decoding to encourage well-formed invocation-end pairs.
@@ 3.3 Memory Formation
Latent Memory Former. Distinct from many existing paradigms [44, 26, 70], we internalize the latent vision memory into lightweight formers, preserving the general abilities of base VLMs and ensuring the compatibility of our paradigm. We initialize two lightweight LoRA adapters, which are respectively designated as the short-term memory former $\mathcal{F}_{s}$ and long-term memory former $\mathcal{F}_{l}$, attached to the vision encoder and the final language model of the VLM, without directly tampering with the core parameters. More precisely, we first append the generated memory query $\mathbf{Q}$ along with a set of learnable memory tokens after the corresponding target token sequence X. Then we process it by short-term or long-term memory former, which contextualizes and embeds the latent memory information:
@@ 3.4 Training Recipe
We design a two-stage training procedure based on GRPO [43], whose optimization objectives are to optimize the effective formation and invocation of latent memory. The first stage enhances the utility of memory, while the second stage maximizes the reward of each invocation, thereby accelerating the convergence of different components steadily. More detailed algorithms and implementations are present in Appendix 7.2 and 8.3.
@@ 4 Experiments
Star Vet Bench Vista Vista -Math Bench Trust
Vanilla [4] 62.6 66.0 54.0 55.4 57.4 59.3 56.0 43.5 67.8 18.9 46.6 52.3 64.8 55.4 57.7 54.5
SFT 64.7 67.5 56.8 54.5 58.7 60.3 57.7 46.1 69.5 22.8 49.0 53.6 67.0 59.1 59.9 56.5
Visual-RFT [35] 65.6 70.5 59.1 58.0 62.9 63.6 62.4 51.7 71.6 26.5 53.0 55.8 70.7 63.2 63.2 59.8
VLM-R1 [44] 66.3 73.0 59.4 60.6 63.8 64.6 63.4 53.0 75.9 34.6 56.7 54.2 69.9 61.7 61.9 61.3
Vision-R1 [26] 67.1 71.7 60.2 60.8 64.0 65.0 63.2 53.9 77.2 38.7 58.2 56.4 72.6 63.6 64.2 62.5
PAPO [66] 64.2 69.8 57.9 53.3 56.7 60.4 61.2 52.5 73.3 34.8 55.5 50.3 67.7 56.5 58.2 58.2
Sketchpad [24] 62.1 64.5 57.0 54.9 52.8 58.3 57.9 47.4 68.4 24.6 49.6 52.1 66.2 57.2 58.5 55.4
GRIT [13] 65.8 67.8 57.9 52.5 51.0 59.0 59.4 51.6 68.1 22.4 50.4 53.7 67.3 60.1 60.4 56.5
PixelReasoner [48] 65.3 67.1 58.7 56.8 60.5 61.7 58.9 49.3 69.6 25.9 50.9 55.9 69.9 61.5 62.4 58.3
DeepEyes [87] 66.4 70.5 60.3 60.4 63.0 64.1 60.3 49.1 70.8 31.5 52.9 57.4 72.6 64.6 64.9 60.5
OpenThinkImg [49] 66.0 71.6 60.8 59.2 61.7 63.9 61.4 52.8 73.0 28.0 53.8 54.9 74.0 64.3 64.4 60.6
Scaffold [28] 63.9 67.0 58.5 52.5 52.9 59.0 58.1 51.0 64.7 21.0 48.7 54.8 68.5 60.6 61.3 56.1
ICoT [16] 65.6 67.9 60.5 54.3 57.0 61.1 58.6 49.8 76.7 30.8 54.0 57.0 69.1 62.0 62.7 59.1
MINT-CoT [9] 66.2 69.5 57.3 55.4 58.9 61.5 57.7 51.5 77.4 39.2 56.5 56.7 71.4 60.8 63.0 60.2
VPT [75] 64.2 70.8 59.0 58.6 63.5 63.2 59.1 53.0 72.3 34.7 54.8 52.3 64.7 61.4 59.5 59.5
Mirage [70] 64.5 71.8 56.1 56.3 59.0 61.5 59.4 50.6 70.3 35.4 53.9 50.9 66.1 60.3 59.1 58.4